HETAV PATEL

Grade 11,

A stone is projected with speed of 5 m/s at an angle of 60°with the horizontal, the speed of stone at highest point of trajectory is ?

A stone is projected with speed of 5 m/s at an angle of 60°with the horizontal, the speed of stone at highest point of trajectory is ?

Grade:12th pass

1 Answers

Mahima Kanawat
6 years ago
Here , U = 5 m/s and angle of projection = 60° At highest point Uy=0 Ux=ucostheta = 5 × cos60=5×0.5=2.5 m/s